OpenStudy (anonymous):

Find an equation of a line perpendicular to y=3x+3 and passing through the point (5, 3).

OpenStudy (anonymous):

a perpendicular line will have the opposite sign and be the reciprical of the current slope of 3, which we have now, so, the slope of the perpendicular line will be as follows: y=mx+b y=3x+3 so m = 3, now take the opposite sign and do the reciprical 3 becomes -3 beocmes -1/3 y=-(1/3)x+b <-so the equation of the new line 3=-(1/3)5+b <- plug in the point (5,3) into the new equation, so we can find 'b' b=-9/5 <- solution to 'b' y=-(1/3)x-(9/5) <-solution to 'an equaiont of a line perpendicular'